Report from Saturday 5th and Sunday 6th July.
By Matt Cambourne, Jon Tooke and Terry Deacon.
A fairly quiet day on shed today as we had D7017 down at Bishops Lydeard for an underframe
examination prior to it heading off to the Gloucestershire and Warwickshire Railway at the end of the
On shed Leroy and Gordon were busy progressing some of the jobs on D1010 in the yard.
John repaired a fault that had occurred recently with Class 47 D1661, whereby a horn valve was not
seating properly when the air pressure built up to a reasonable level. This resulted in the horn constantly
sounding! Now this is fixed the loco can be used on the remaining running days this year.
Hymek D7018 continues to be progressed. Matt fitted a new engine hour counter today including a
button which enables the system to display the reading with the engine shut down. Air pressure and
vacuum gauges are currently being calibrated and repaired where required so we have a good full
working set for when the loco is up and running.
David was busy working on the cooler group for D9518 today, this is the first component being tackled in
the overhaul and has been done to a very high standard.
Meanwhile at Bishops Lydeard as Jon Tooke reports, much activity was taking place on D7017:-
Today, Saturday, in readiness for the locos planned visit to the GWR at Toddington, a team from the
DEPG of Graham, Martin , Bob C, Jon T and Simon took the Hymek down to BL and placed the loco on
one of the pit roads in the compound for a thorough inspection, which if you look at old photos
looks a bit like Horton Road depot in Gloucester from years gone by!
When unloaded at GWR, another inspection will take place to make sure that nothing untoward occurred
during the move by road transport.
The locomotive is pictured at Bishops Lydeard below.
On Sunday Terry and the Class 14 crew were down at Williton preparing D9526 for its forthcoming visit
to the East Lancs Class 14 at 50 event. Pictures below from Terry.
Below that are some further pictures from Jon showing the various locos on the depot as we undertook a
shunt to get D7017 out to go to Bishops Lydeard.